OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4)and (2, -2)? A.) negative three halves B.) negative two thirds C.) three halves D.) two thirds

OpenStudy (anonymous):

can u calculate the slope between the two points?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and what do you have?

OpenStudy (anonymous):

(subtract the y's) over (subtract the x's)

OpenStudy (anonymous):

subtract the y's: 4 - (-2) = subtract the x's: -2 - (2) =

OpenStudy (anonymous):

6 & -4?

OpenStudy (anonymous):

ok... but as a fraction...

OpenStudy (anonymous):

i mean. 6/-4

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\huge \frac{6}{-4}= \]

OpenStudy (anonymous):

ok... and can you reduce that fraction?

OpenStudy (anonymous):